simplify the expression

sin x cos x
--------------
tan x
1
Expert's answer
2012-12-20T08:59:23-0500
tan(x) = sin(x) / cos (x)
hence
sin(x)cos(x) / tan(x)= sin(x)*cos(x)*cos(x) / sin(x) = cos(x)*cos(x) = (cos(x))^2